OpenStudy (anonymous):

ok can somebody plz help me and walk through with this step by step . The lengths of two sides of a right triangle are given. Find the length of the third side. Round to the nearest tenth if necessary. legs:28 in. and 15in.

OpenStudy (anonymous):

a^2+b^2=c^2 28^2+15^2=

OpenStudy (anonymous):

c^2=1009 square root of 1009 31.76

OpenStudy (anonymous):

ok thanks not sure if i under stand .

OpenStudy (anonymous):

a^2+b^2=c^2 is the formula a and b stands for the two legs and c is the hypotenuse

OpenStudy (anonymous):

yea i know that

OpenStudy (anonymous):

so what do i do add them to numbers